The mass of a box measured by a grocer.s balance is 2.300 kg. Two gold pieces of
masses 20.15 g and 20.17 g are added to the box. What is
(a) the total mass of the
box,
(b) the difference in the masses of the pieces to correct significant figures ?
Grade:11

1 Answers

Kevin Nash
10 years ago

Mass of grocer’s box = 2.300 kg

Mass of gold piece I = 20.15g = 0.02015 kg

Mass of gold piece II = 20.17 g = 0.02017 kg

(a) Total mass of the box = 2.3 + 0.02015 + 0.02017 = 2.34032 kg

In addition, the final result should retain as many decimal places as there are in the number with the least decimal places. Hence, the total mass of the box is 2.3 kg.

Ans b 

(b) Difference in masses = 20.17 – 20.15 = 0.02 g

In subtraction, the final result should retain as many decimal places as there are in the number with the least decimal places.